Fermi National Laboratory


Accelerator Update for July 13, 2005 - July 15, 2005

Wednesday July 13
The day shift began with Operations monitoring Tevatron (TeV) store 4269 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following colliding, neutrino, and fixed target experiments taking beam: the Collider Detector at Fermilab (CDF), D-Zero (D0), MiniBooNE, NuMI, and the Meson MCenter experiment.

A Pbar expert began a study period at 8:45 AM.

TeV experts began an end-of-store study around 9:30 AM.

Operations terminated store 4269 at 10:57 AM. TeV experts began a proton only study.

Pbar studies ended at 1:14 PM.

At 2:04 PM, the Tevatron Electron Lens (TEL) tripped off when the TeV CUB LCW went down. It came back up when the CUB work finished and the TeV LCW system was brought back on.

Operations loaded a 36x0 study-store in the TeV at 4:05 PM.

Operations loaded a 4x0 study-store in the TeV at 5:18 PM.

Operations put the TeV through a dry squeeze at 5:30 PM.

Operations put the TeV into shot setup at 6:45 PM.

At 6:53 PM, Operations reported that the Main Injector (MI) beam permit had tripped off for the first of four times in eleven minutes. It reset each time.

At 8:01 PM, Operations reported that feeder 37 tripped off and caused a cryo system cold compressor to trip off at CHL. The High Voltage Coordinator investigated the feeder trip.

At 8:07 PM, a Booster RF station (BRF12) tripped off for the first of three times in one hour on a ground fault indication. It reset each time.

The Cryo Coordinator reported at 8:36 PM that the cryo system was too unstable for the TeV to ramp. The shot setup was put on hold until cryo technicians can get another compressor online.

The cryo tech at CHL reported at 9:24 PM that it was safe to continue with shot setup.

Operations established combination store 4271 at 10:31 PM with an initial luminosity of 91.72E30. Pbar resumed stacking.

At 11:52 PM, Operations reported that the store 4271ís lifetime had been decaying. TeV experts investigated.

Thursday July 14
The midnight shift began with Operations monitoring store 4271, with Pbar stacking, and with CDF, D0, MiniBooNE, NuMI, and MCenter taking beam.

At 12:17 AM, the TeV suffered a multi-house quench.

Operations prepared D0 for an access into their collision hall beginning at 2 AM.

Operations put the Recycler into shot setup at 2:29 AM to stash antiprotons. The stash ended and stacking resumed at 3:43 AM.

At 5:58 AM, Operations reported that the Pbar antiproton target had stopped rotating. Operations will inform Pbar experts when they come in this morning.

D0 came out of access at 6:05 AM.

Operations put the TeV through a dry squeeze at 6:40 AM.

At 7 AM, TeV experts began investigating the problems that had caused the store to quench.

Operations loaded a 36x0 study-store at 8:18 AM and then another one at 9:10 AM.

Pbar resumed stacking at 10:23 AM.

Operations began loading another 36x0 study-store for the TeV experts at 11:55 AM.

Operations terminated the 36x0 study-store at 2:22 PM.

At 4:33 PM, Operations reported that NuMI had been tripping off due to a kicker failing to fire. Expert investigated.

Operations loaded another 36x0 study-store at 8:48 PM.

Operations put the TeV into shot setup at 9:28 PM.

Between 11 PM and midnight, Operations reported various problems transferring antiprotons from the Recycler and Pbar to the TeV.

Friday July 15
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established combination store 4276 at 12:55 AM with an initial luminosity of 78.78E30. Pbar resumed stacking.

MCenter resumed taking beam at 12:59 AM.

The Recycler began stashing antiprotons at 6:35 AM.

The stash ended and Pbar resumed stacking at 7:42 AM.

MCenter resumed taking beam at 7:50 AM.

The Plan for Friday and the Weekend
The plans for today are to monitor the store and conduct machine studies. The rest of the weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/13/2005   email Fermilab